A particle of charge q and mass m starts its motion from origin with u=a i.A uniform B=b/4 i + (sin 60)/2 j exists everywhere in space.Find the component of v in y dir. when z coord. of it becomes max.

Sir, if I find the magnetic force,Its component is 0 in Y direction.With no force acting in Y direction ,how come its velocity changes??

Plz help me!!!


Grade:upto college level

1 Answers

ROSHAN MUJEEB
3 years ago


Letvelocityisgivenby:V=aiˆ+ujˆwhereuisthecomponentofvelocityinydirectionB=b4iˆ+3√4jˆforceisgivenby:f=q(v×B)f=q(aiˆ+ujˆ)×(b4iˆ+3√4jˆ)f=(3√a4−ub/4)kˆsinceforceisinzdierection,andifzcomponentoftheforcemustbemaximumthenumusthaveleastpossibleval;uewhichiszero.thereforeycomponentovelocitymustbezero.